L=3mH . How to find the Reactive power given the following circuit :
5UF + Vout W = 4.635 rad/s Vin = 2220° + To 40.22 Vin R2 20-2 ZRA = 402 ZR2 = 2012 ZL = jw (3e - 3) Zc = 1 jw(5e-6) jwc
Answer #1

SUF L=3mt riven ] در مرا T! Ripo 2010 Wo 9.635 rad see steps: 1) c = wc 9635x5x156 = 20-751 1) XL= احيا 9635 X 3x103 XL = 28-(བྱ༢ 5 ste PH ke at junction Applying Irt12 0.488 127.941 of 0.605152-724 0-8398/17-801 A steps I 20.8398 [17.801 A 22lova i